OpenStudy (dls):

The kinetic energy acquired by a mass m in traveling distance d,starting from rest,under the action of a constant force is directly propotional to? A)m B)m^0 C)m^1/2 D)m^-1/2

OpenStudy (anonymous):

really!? work done = change in kinetic energy! :-/

OpenStudy (anonymous):

Mashy, despite of how mean that answer was, it is correct.

OpenStudy (anonymous):

he is my bro.. so don't worry :D

OpenStudy (anonymous):

Lmao okay

OpenStudy (anonymous):

and he is appearing for IIT :D.. what a joke :D

OpenStudy (dls):

-.- sorry it was my net yesterday..couldnt post options.. A)m B)m^0 C)m^1/2 D)m^-1/2

OpenStudy (dls):

I thought of it like this.. constant force so acceleration = 0 v=constant so directly propotional to m? but its wrong..

OpenStudy (dls):

@experimentX,,okay something serious

OpenStudy (experimentx):

work done...

OpenStudy (dls):

options are A)m B)m^0 C)m^1/2 D)m^-1/2

OpenStudy (dls):

I got A but its wrong

OpenStudy (experimentx):

most likely B ...

OpenStudy (dls):

y?

OpenStudy (experimentx):

there is constant force, ... and the 'd' is fixed, not matter what your mass is ... it will acquire the same KE.

OpenStudy (dls):

numerically it should be m but?

OpenStudy (experimentx):

you are applying same force to all ... along the same distance, ... it doesn't matter if your object is sum or an atom, the work done must be same ... which is equal to it's KE

OpenStudy (shubhamsrg):

const force, so accn 0 ,, wow! :O

OpenStudy (yrelhan4):

so you want a numerical answer right? distance travelled=d=v^2/2a where a-->acceleration by manipulating s=K/F . so distance traveled is independent of mass. and if you think K=1/2mv^2, here v is not constant, so you cannot use this.

OpenStudy (shubhamsrg):

NO

OpenStudy (dls):

oops

OpenStudy (experimentx):

I think constant force will produce constant acceleration.

OpenStudy (dls):

if force is constant then how will there be difference in acceleration o.O

OpenStudy (dls):

yes const.acc

OpenStudy (experimentx):

imagine a falling object ... the force is always constant, for small distance.

OpenStudy (dls):

that happens in SHM ! :O

OpenStudy (dls):

\[F = -kx\]

OpenStudy (anonymous):

omg work done = change in kinetic energy work done = Fd F= ma hence work done = mad (lol) since initial k.E = 0.. change in K.E = Final K.E - 0 = Final K.E hence Final K.E = mad or Final k.E directly proportional to m !!

OpenStudy (dls):

i think i said the same? but its B) and I got it too.. also you did it wrong

OpenStudy (anonymous):

the kinetic energy acquired IS proportional to mass.. think logically your force is constant.. the distance for which you are putting the force is ALSO constant so imagine throwing a BIGGGG boulder.. and imagine throwing a small rock!! in both case.. the force you put on them is fairly same? .. cause it depends on your strength? and your swing action is more or less the same?? hence the distance is the saem?? now you are saying both the boulder and the small rock gain the same speed K.E?

OpenStudy (anonymous):

WAIT WAIT WAIT...i got i got it! :D

OpenStudy (shubhamsrg):

the force is constant, independent of the mass! -_-

OpenStudy (anonymous):

they get same K.E... but not same speed!!

OpenStudy (shubhamsrg):

in this ques.

OpenStudy (anonymous):

yes yes.. correct.. !! :P independent of mass!!

OpenStudy (anonymous):

i get it.. i get it..!! hope DLS gets it too!?

OpenStudy (shubhamsrg):

Its not possible that DLS doesn't get it! ^_^ He's a champ! \m/

OpenStudy (dls):

\[\LARGE WD=FORCE \times DISPLACEMENT\] force=constant displacement=constant WD=constant

OpenStudy (anonymous):

wow shubh.. that was INTENSE sarcasm.. of a very high level :D

OpenStudy (dls):

..

OpenStudy (shubhamsrg):

DLS (Dynamic light scattering (also known as photon correlation spectroscopy or quasi-elastic light scattering))

OpenStudy (anonymous):

now he is just braggin :P

OpenStudy (shubhamsrg):

That wasn't sarcasm! ^_^

OpenStudy (dls):

i got confused again

OpenStudy (anonymous):

why confused? just like you said.. work done is constant.. hence k.e is also constant!

OpenStudy (dls):

i mean in the other question :/

OpenStudy (anonymous):

So what's the net result..I am literally confused

OpenStudy (anonymous):

the result is that.. K.E is independent of Mass in this case

OpenStudy (anonymous):

How does this fits in , in a real life scenario?

OpenStudy (anonymous):

it does fit in.. the real life example i gave.. it works for that too!